An airplane while taking off moves with a constant
acceleration over a runway of 600 m in 10 seconds.
Determine the velocity with which it takes off. Also,
determine its constant acceleration.
1
Expert's answer
2019-12-12T06:31:50-0500

"s=\frac{1}{2}at^2"

"a=12\frac{m}{sec^2}"

"v=at"

"v=120\frac{m}{sec}"